OpenStudy (anonymous):

What is the solution to the equation? 6 + 2s – 8s = 18 A. s = –4 B. s = –2 C. s = 2 D. s = 4

OpenStudy (anonymous):

first try plugging in the numbers for s

OpenStudy (anonymous):

Here: Simplify 6 + 2s + -8s = 18 Combine like terms: 2s + -8s = -6s 6 + -6s = 18 Solve 6 + -6s = 18 Solve for variable 's'. Move all terms containing s to the left, all other terms to the right. Add '-6' to each side of the equation. 6 + -6 + -6s = 18 + -6 Combine like terms: 6 + -6 = 0 0 + -6s = 18 + -6 -6s = 18 + -6 Combine like terms: 18 + -6 = 12 -6s = 12 Divide each side by '-6'. s = -2 Simplify s = -2 Another way to do it, is to put in one of the answer choices in the equation one at a time in place of 's' - to see which one makes the equation true!

OpenStudy (anonymous):

6+2S-8S=18 -6 from both sides 2S-8S=12 add like terms (2S-8S) -6S=12 Divide -6 from both sides S=-2

OpenStudy (anonymous):

Thank you @mangorox you seem to make me understand everything very well thank you !

OpenStudy (anonymous):

If I could I would medal all your guys responses! ^_^

OpenStudy (anonymous):

lets try the first one: 6+2(-4)-8(-4)=18 so 2*-4=-8 add 6 to that: 6+(-8)=2 now for 8(-4)= -32 oops nvm @mangorox got it xD